OpenStudy (anonymous):

Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. A(n) = -6 + (n-1) (1/5)

OpenStudy (jhannybean):

\[\large a_{n}= a_{1}+(n-1)d\] This is the formula for the Arithmetic Progression.\[\large a_{n}=-6+(n-1)(\frac{1}{5})\]your formula \[\large a_{1}=-6+(1-1)(\frac{1}{5})\]\[\large a_{4}=-6+(4-1)(\frac{1}{5})\]\[\large a_{10}=-6+(10-1)(\frac{1}{5})\] Simplifying these you should be able to get your first, fourth and 10th terms :) I hope this helps!